WARNING: This web is not used anymore. Please use PDBService.MinuteS8February07 instead!
 

Minutes 8 February 07

  • Phone:
    • TRIUMF: Jim Li
    • GridKA: Doris
    • FNAL: Eric
    • ATLAS: Shasha
    • BNL: Iris, Carlos
    • NDGF: Olli
    • CNAF: Barbara

  • CERN:
    • 3D: Dirk, Luis, Eva
    • ATLAS: Gancho, Florbela
    • LHCb: Marco

  • General announcements:
    • GDB meeting:
      • presentations by Maria (technical side) and Sasha (experiment side)
      • backup policy has been approved
      • requests from ATLAS: 3rd node for the RAC installations - discussion about the real need, the sites should be prepared and check what is needed to meet the requirements
    • PIC has hired a new dba - Luis Diaz, he will come to CERN for training
    • Workshop @SARA
      • preparation for the dba meeting ongoing
      • agenda still to be completed
      • Dirk will propose some topics
      • dbas should prepare the issues to be discussed
      • Jamie Shiers will join the workshop to talk about the LCG milestones and experience

  • Sites status
    • OEM installation to be completed
    • The sites must contact their ROC
    • BNL (Iris):
      • confirmed that the network problems are caused by the firewall
      • the oracle cluster will be moved outside the firewall
      • oem agents are down due to this work
      • ROC already contacted
      • metalink account: done
    • CNAF (Barbara):
      • problems with the new storage: there will be an intervention by the storage experts next monday
      • in the meantime, LHCb and ATLAS databases are available using the old storage
      • ROC already contacted
      • metalink account: actually using previous CNAF account. They should get new account for the SR related to the 3d project
      • patches will be applied when the storage problem is fixed
    • GridKA (Doris):
      • question about backups: how to estimate the space?
      • patches have been applied successfuly
      • oem agents up and running
      • Doris does not know whether ROC was contacted or not, she will check
      • metalink account: done
    • NDGF (Olli):
      • db using single instance up and running
      • patches have been applied
      • it is necessary to determine if the single instance database will be part of the tests - check if streams + oem can maintain the throughput
      • ROC already contacted
      • metalink account: done
    • RAL: still problems with the email list. One instance of LHCb cluster is down (problems during the patches apply)
    • SARA: pathes have been applied, oem agents up and running
    • Triumf (Jim Li):
      • iperf tests show very low network throughput
      • Jim Li should check if there is a firewall between CERN and Triumf
      • more tests will be run ( suggested by CERN network experts)
      • oem agent up and running
      • ROC to be contacted

  • Experiments status:
    • ATLAS:
      • downstream configuration is ready
      • first tests with RAL as destination were successful (Eva)
      • GriKA and SARA joined: test during 27 hours was ok
      • IN2P3 and CNAF will be joined soon
      • IN2P3 will also be joined to the direct setup for Richard's tests
      • Michigan setup is ready - we should get a plan about their tests
      • move the Streams configuration to production to be discussed
    • LHCb:
      • CNAF to be added: data can be discarded
      • downstream capture configuration will be done tomorrow
      • discussion about the LFC: inside the LHCb cluster or will be a separate instance
      • LFC sw is being upgraded to allow bulk operations
      • schema changes will be done at the master database
      • new replicas will be added after the schema changes
      • LHCb will run COOL 2.0 for production

  • CMS Frontier tests:
    • working with the online people to run the tests
    • 2 squid nodes have been requested at Tier0 (FIO)
    • as Luis is leaving, it is necessary to agree between CMS and FIO how to proceed with notifications, reboots, ...
Edit | Attach | Watch | Print version | History: r1 | Backlinks | Raw View | WYSIWYG | More topic actions
Topic revision: r1 - 2007-02-23 - EvaDafonte
 
    • Cern Search Icon Cern Search
    • TWiki Search Icon TWiki Search
    • Google Search Icon Google Search

    PSSGroup All webs login

This site is powered by the TWiki collaboration platform Powered by PerlCopyright & 2008-2020 by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
or Ideas, requests, problems regarding TWiki? use Discourse or Send feedback